OpenStudy (anonymous):

Miguel leaves by automobil at 9:00 A.M. and stops for repairs at 9:20 A.M. If the distance covered was 18 miles, what was the average velocity, in miles per hour for this part of the trip?

ganeshie8 (ganeshie8):

20 minutes = 1/3 hour

ganeshie8 (ganeshie8):

\(\large \text{Average velocity} = \frac{\text{displacement}}{\text{time}}\)

ganeshie8 (ganeshie8):

\(\large \text{Average velocity} = \frac{\text{18 miles }}{\text{1/3 hour}}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

Oh thanks :)

ganeshie8 (ganeshie8):

np :)